OpenStudy (anonymous):

WILL FAN AND MEDAL A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. 5x + y = −6 C. (y – 4) = −5(x + 2) D. (y + 4) = −5(x − 2)

OpenStudy (anonymous):

A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. 5x + y = −6 C. (y – 4) = −5(x + 2) D. (y + 4) = −5(x − 2)